OpenStudy (anonymous):

Choose the correct simplification of 2a0.

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

what is a number raised to 0?

OpenStudy (anonymous):

0 is an exponet

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so it would be 0?

terenzreignz (terenzreignz):

Yes. When 0 is the exponent... we have a property... \[\huge a^0 = \color{red}?\]

OpenStudy (anonymous):

0

terenzreignz (terenzreignz):

No... when something is raised to a 0 exponent, it is equal to 1, ok? \[\huge a^0 =\color{blue}1\]

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

So what's \[\huge 2a^0\] based on what you just found out?

OpenStudy (anonymous):

1

terenzreignz (terenzreignz):

NO! \[\huge a^0 = 1\] \[\huge \color{green}2a^0 = \color{red}?\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

or 2a?

terenzreignz (terenzreignz):

Good. Glad we straightened that out :)

OpenStudy (anonymous):

so its 2a

terenzreignz (terenzreignz):

Well, you decide... which is it?

OpenStudy (anonymous):

2a

terenzreignz (terenzreignz):

I thought I made it clear to you that \[\huge \color{red}{a^0 = 1}\]? If it's not clear to you yet, digest it~

OpenStudy (anonymous):

oooo so the a goes away and its just 2

terenzreignz (terenzreignz):

Yes. Precisely. Remember this, and remember this well... ANYTHING when raised to a 0 exponent just becomes 1.

OpenStudy (anonymous):

okay thnx

terenzreignz (terenzreignz):

Okay. I have to go now. ---------------------------------------- Terence out